Consider two individuals, Bill and Steve. Both can make cars and computers. Their production levels are outlined below:

In one hour of time Bill can make: 7 Cars -or- 35 Computers.
In one hour of time Steve can make: 4 Cars -or- 12 Computers.

1. What is Bill's opportunity cost of making one (1) Car?
What is Steve's opportunity cost of making one (1) Car?

2. What is Bill's opportunity cost of making one (1) Computer?
What is Steve's opportunity cost of making one (1) Computer?

3. a. Who has the absolute advantage in Car production? Explain.
b. Who has the absolute advantage in Computer production? Explain.

4 a. Who has comparative advantage in Car production? Explain.
b. Who has comparative advantage in Computer production? Explain.

5. In the event the two decide to specialize and trade away what they are better in making, who will produce what? What will the terms of trade be? Who is going to benefit and by how much? Please be specific.
Answer #1

1) Bill can make 7 cars or 35 computers in 1 hour.

Bill's opportunity cost of making 1 car = 35/7 = 5 computers.

Steve can make 4 cars or 12 computers in 1 hour.

Steve's opportunity cost of making 1 car =12/4 =3 computers.

2) Bill's opportunity cost of 1 computer = 1/5 cars

Steve's opportunity cost of 1 computer = 1/3 cars

3) a) Bill has absolute advantage in car production as he can produce more of it in 1 hour.

b) Bill has absolute advantage in computer production as he can produce more of it in 1 hour.

4) a) To make 1 car, the opportunity cost of Bill and Steve is 5 computers and 3 computers respectively. As Steve sacrifices less computers to produce cars, Steve has comparative advantage in car production.

b) To make 1 computer, the opportunity cost of Bill and Steve is 1/5 cars and 1/3 cars respectively. As Bill sacrifices less cars to produce computers, Bill has comparative advantage in computer production.

5) Bill will produce 35 computers and Steve will produce 12 computers. Terms of trade is 35/12.Bill can give 12 computers to Steve and Steve can give 4 cars to Bill in exchange for computers. Bill will benefit by having 35 - 12 = 23 computers and 4 cars. Steve's benefit will depend on how many computers Bill gives to Steve.